# Trustworthy Transfer Learning: A Survey

### v1

@@ 3.2 Non-IID Transferability
The IID assumption is often violated in real scenarios, e.g., connected nodes in graphs (?), word occurrence in texts (?), temporal observations in time series (?), etc. To bridge this gap, non-IID transferability explores the knowledge transfer across domains, assuming that samples within each domain can be interdependent.
@@
Figure 4: Illustration of distribution shifts in characterizing graph transferability. The node distribution within the graph can be represented by $\mathrm{P}(X,G,Y)$ where $X,G,Y$ denote the input node attributes, topology structure, and output class labels, respectively. The color of nodes indicates the class labels $Y$ (blue or green).
3.2.1 Transferability on Graph Data
@@
Graph data is being generated across a variety of application domains, ranging from bioinformatics (?) to e-commerce (?), from protein-protein interaction prediction (?) to social network analysis (?). To capture the complex structure of graph data, graph neural networks (GNNs) (?, ?) have been introduced to encode the nodes within the graphs into low-dimensional vector representations. It is shown (?) that there are two major learning paradigms for GNNs: spectral-based (?, ?) and spatial-based GNNs (?, ?, ?). Recently, the transferability of spectral and spatial GNNs has been studied by exploring whether GNNs are transferable across graphs of varying sizes and topologies (?, ?). In this survey, we focus on understanding the transferability of GNNs in node-level graph learning tasks. Note that the key challenge to theoretically understand the transferability of GNNs is to measure the distribution shifts of two graphs. As illustrated in Figure 4, the distribution shifts between source and target graphs are generally induced by the joint probabilities $\mathrm{P}_{S}(X,G,Y)$ and $\mathrm{P}_{T}(X,G,Y)$, where $X,G,Y$ denote the input node attributes, topology structure, and output class labels, respectively. Specifically, the transferability of spectral GNNs leverages the graph limits, e.g., graphon (?, ?) or graphop (?), to determine if two graphs represent the same underlying structure as the number of nodes goes to infinity. In contrast, the transferability of spatial GNNs typically relies on the empirical distribution differences of node representations in a latent embedding space learned by GNNs (?, ?).
@@
Spectral GNNs define the graph convolutions in the spectral domain using the graph Fourier transform from the perspective of graph signal processing (?, ?). Recent efforts (?, ?, ?, ?, ?) have been dedicated to understanding the transferability of spectral GNNs by answering the following question: Can spectral GNNs trained on a source graph perform well on a target graph of different sizes? This question is also known as size generalization (?, ?). The intuition behind the transferability of spectral GNNs is that if two graphs represent the same underlying phenomenon, their GNN outputs will be similar. Thus, the transferability of spectral GNNs can be derived from various aspects, including generic topological space (?), graphon (?), graphop (?), and $k$-hop ego-graph (?). To be more specific, ? (?, ?) study the transferability of spectral graph filters on different discretizations of the same underlying continuous topological space. Later, graphon theory (?) is used to analyze the transferability of spectral GNNs. Formally, a graphon is defined by a bounded symmetric kernel and can be viewed as a graph with an uncountable number of nodes. In particular, ? (?) leverage graphon to study the asymptotic behavior of GNNs (?), showing that GNNs converge to graphon neural networks (WNNs) as the number of nodes increases to infinity. This convergence implies that under mild assumptions, GNNs are transferable across graphs with performance guarantees if both graphs are drawn from the same graphon (?, ?, ?). Following this observation, recent works (?, ?) further demonstrate the transferability of the gradients of spectral-based GNNs across graphs under similar conditions. Furthermore, using the graphop operator (?) which is a generalization of graphon, ? (?) extend the transferability analysis of GNNs to both dense and sparse graphs. Besides, assuming that the $k$-hop ego-graphs are independent and identically drawn, ? (?) derive the transferability of a well-designed GNN based on the differences of $k$-hop ego-graph Laplacian across graphs.
@@
Spatial GNNs generally follow a recursive message-passing scheme (?), where each node updates its feature vector by aggregating the message from its local neighborhood. As discussed in (?, ?), the marginal distribution shifts $\mathrm{P}(X,G)$ between source and target graph domains can be induced by graph structure and individual node attributes (see Figure 4(a)-(c)). Notably, three frameworks have been developed to enhance the transferability of spatial GNNs: invariant node representation (?, ?, ?), structure reweighting (?, ?), and graph Gaussian process (?).
@@
    Invariant Node Representation: Inspired by the domain adaptation theory (?), it is theoretically shown (?, ?) that the target error can be bounded in terms of the source error and the graph domain discrepancy. The crucial idea of invariant node representation learning is to explicitly minimize the graph domain discrepancy in a latent feature space, thereby enhancing the transferability of spatial GNNs. For example, AdaGCN (?) and UDA-GCN (?) leverage a domain discriminator to learn the domain-invariant node representation learned by the output layer of GNNs. Inspired by the connection between spatial GNNs and Weisfeiler-Lehman graph kernels (?, ?), GRADE (?) is proposed based on a graph subtree discrepancy measuring the subtree representation induced distribution shifts across graphs. More recently, SpecReg (?) and A2GNN (?) further discuss the impact of spectral regularization and asymmetric model architectures on the transferability of GNNs, respectively.
@@
    Structure Reweighting: It is noticed (?) that invariant node representation might lead to sub-optimal solutions under conditional structure shifts. To solve this problem, StruRW (?) and Pair-Align (?) are proposed to reweigh the edges of the source graph based on the label-oriented node connections of source and target graphs.
@@
    Graph Gaussian Process: Spatial GNNs are equivalent to graph Gaussian processes in the limit as the width of graph neural layers approaches infinity (?). Based on this observation, GraphGP (?) is derived from a graph structure-aware neural network in the limit on the layer width, in order to characterize the relationships between nodes across different graph domains. The generalization analysis of GraphGP further reveals the positive correlation between knowledge transferability and graph domain similarity.
3.2.2 Transferability on Textual Data
@@
Transfer learning has been widely studied in various natural language processing (NLP) tasks, e.g., text classification (?), question answering (?), neural machine translation (?), etc. A key challenge in understanding textual transferability is the non-IID nature of words/tokens, as they might co-occur within sequences or documents. Thus, recent theoretical analyses of textual transferability often consider an alternative assumption (?), i.e., sequences or documents are independently drawn from the same distribution. This assumption enables theoretically deriving the transferability and generalization of transfer learning in sequence-level and document-level NLP tasks. Take multilingual machine translation as an example, the goal is to train a single neural machine translation model to translate between multiple source and target languages (?). To achieve this, language-invariant representation learning has been introduced to align the sentence distributions of different languages within a shared latent space (?). Nevertheless, ? (?) theoretically analyze the fundamental limits of language-invariant representation learning by deriving a lower bound (w.r.t. marginal sentence distributions from different languages) on the translation error in the many-to-many language translation setting.
@@
More recently, large language models (LLMs) have revolutionized the field of NLP (?, ?, ?, ?, ?). The transferability of LLMs has been studied, as fine-tuning LLMs on downstream tasks has become the de facto learning paradigm. However, it is computationally expensive and resource-intensive to fine-tune the entire LLM model weights with billions of parameters via gradient-based optimization (?). To solve this problem, parameter-efficient fine-tuning (PEFT) has been investigated from the perspectives of model tuning (?, ?, ?) and prompt tuning (?, ?, ?). The goal is to adapt LLMs to various downstream tasks by adjusting as few parameters as possible.
@@
Model tuning based approaches explore model architectures or parameters of LLMs for parameter-efficient fine-tuning (?). A variety of parameter-efficient model tuning frameworks have been proposed, including adapters (?), low-rank decomposition (?, ?), and selective masking (?, ?). The key ideas behind these frameworks are illustrated in Figure 5. In general, these approaches aim to update only a few parameters by inserting new trainable modules, adding low-rank parameter matrices, or modifying specific parameters (e.g., bias terms).
@@
    Adapters: Adapter-based approaches add new learnable modules with a small number of parameters to LLMs, e.g., maximizing the likelihood $p(y|x;\theta_{\mathrm{adapter}}\circ\theta_{\mathrm{LLM}})$ with added modules $\theta_{\mathrm{adapter}}$. Initially, inspired by visual adapter modules (?), ? (?) study the adapter-based fine-tuning mechanism in NLP tasks. This method inserts two adapters sequentially within each Transformer block (?): one following the self-attention layer and another after the feed-forward layer. Nevertheless, follow-up research (?, ?) demonstrates that inserting a single adapter after the feed-forward layer can achieve competitive performance while adding fewer parameters. Furthermore, by highlighting the connections between (model-based) adapters and (prompt-based) Prefix Tuning (?), ? (?) introduce a family of parallel adapters that directly condition the adapters at different Transformer layers directly on the input text. On top of adapters, AdapterDrop (?) and CoDA (?) improve both fine-tuning and inference efficiency, by removing adapters from lower Transformer layers and querying only a small subset of input tokens against the pre-trained LLMs, respectively.
    Figure 5: Illustration of parameter-efficient fine-tuning with (b) Adapters (?) where new modules are inserted, (c) LoRA (?) where low-rank parameter matrices are added, and (d) BitFit (?) where only bias terms will be updated.
@@
    Low-rank Decomposition: Low-rank decomposition injects trainable low-rank decomposition matrices into pre-trained model parameters without changing the model architectures, e.g., maximizing the likelihood $p(y|x;\theta_{\mathrm{LLM}}+\Delta\theta_{\mathrm{LLM}})$ with low-rank $\Delta\theta_{\mathrm{LLM}}$. This line of research is motivated by the phenomenon (?, ?) that pre-trained models tend to have a low intrinsic dimension. Here, the intrinsic dimension indicates the lowest dimensional parameter subspace in which satisfactory fine-tuned accuracy on downstream tasks can be achieved. Moreover, by assuming that the parameter change in LLMs during fine-tuning also has a low ”intrinsic rank”, LoRA (?) is introduced by optimizing low-rank decomposition matrices of the parameter change. Empirically, LoRA has been further improved in various aspects, including rank selection/optimization (?, ?, ?), advanced optimizer (?, ?) etc. Theoretically, the expressiveness and generalization of LoRA have also been analyzed. Notably, ? (?) prove that under mild conditions regarding the rank of LoRA, it can adapt a pre-trained (or randomly initialized) model to approximate any target model of equal or smaller size. ? (?) find that LoRA fine-tuning is approximately equivalent to full fine-tuning in the Neural Tangent Kernel (NTK) regime (?), if $r\geq\Theta(\log n^{t}/\epsilon^{2})$ where $r$ is the rank of LoRA and $\epsilon$ is an approximation tolerance. ? (?) show that LoRA fine-tuning has no spurious local minima in the NTK regime, if $r(r+1)>2Kn^{t}$ where $K$ is the output dimension, and $n^{t}$ is the number of training samples in the downstream target task. Furthermore, the generalization bounds of LoRA fine-tuning are theoretically derived in recent works (?, ?).
@@
    Selective Masking: The crucial idea of selective masking is to update only a small subset of model parameters during fine-tuning, e.g., maximizing the likelihood $p(y|x;\theta_{\mathrm{LLM}}+\Delta\theta_{\mathrm{LLM}})$ with extremely sparse $\Delta\theta_{\mathrm{LLM}}$. Intuitively, it aims to find a binary mask that automatically selects a small subset of parameters for fine-tuning (?). There are three main frameworks for learning this mask. The first one is random masking (?, ?), where all the elements of a mask are sampled independently from a Bernoulli distribution. ? (?) demonstrate the effectiveness of random masking under a larger-than-expected learning rate, by theoretically building the connection between random masking and flat loss landscape. The second approach involves heuristically-motivated masks, such as bias terms (?), layer normalization modules (?), and cross-attention layers (?). The third approach optimizes masks over model parameters from various perspectives, e.g., $L_{0}$-norm penalty (?), Fisher information (?, ?, ?), Lottery Ticket Hypothesis (?, ?), etc.
Unlike model tuning, prompt tuning keeps the pre-trained LLMs fixed and pretends a sequence of virtual token embeddings (referred to as a trainable prompt) to the input text, e.g., maximizing the likelihood $p(y|[x,z];\theta_{\mathrm{LLM}})$ for a labeled sample $(x,y)$ where $\theta_{\mathrm{LLM}}$ denotes LLM model parameters and $z$ represents a prompt. Generally, there are three main frameworks for optimizing newly added prompts: soft (continuous) prompt tuning (?, ?), hard (discrete) prompt tuning (?), and transferable prompt tuning (?).
@@
    Soft Prompt Tuning: The key idea of soft prompt tuning is to represent the virtual prompt as continuous-valued token embeddings. It will update only the continuous-valued embeddings of prompts $z$, either in the input embedding layer (?, ?) or in different layers of the LLMs (?, ?, ?, ?). Theoretically, ? (?) studies the connection between prompt tuning and downstream tasks using an underlying latent variable generative model of text. By assuming that input texts are generated by a Hidden Markov Model (HMM), this work models the downstream task as a function of the posterior distribution of the latent variables. It is then shown that prompt tuning enhances the recovery of the ground-truth labeling function in the downstream classification task. Later, ? (?) further show that a carefully constructed pre-trained Transformer can leverage prompt tuning to approximate any sequence-to-sequence function in a Lipschitz function space. They also analyze the restricted expressiveness of prompt tuning compared to model fine-tuning (e.g., LoRA (?)).
@@
    Hard Prompt Tuning: Although soft prompts can be optimized via gradient-based optimization, ? (?) reveal that the learned embeddings of soft prompts do not correspond to any human-readable tokens, thus lacking semantic interpretations. An alternative solution is hard prompt optimization (?, ?), which aims to find human-readable prompts from a pre-defined vocabulary. Specifically, AutoPrompt (?) greedily selects the optimal token for each location in the prompt based on the gradient of the loss w.r.t. the embeddings over labeled training samples. However, the greedy search strategy can result in disfluent and unnatural prompts. To solve this problem, FluentPrompt (?) and PEZ (?) utilize projected gradient descent optimization to update all the tokens in the prompt. The crucial idea is to project the learned continuous-valued embeddings to their nearest neighbors in a pre-defined discrete token space, and then use the mapped tokens to calculate the gradient of the loss. Besides, (?, ?) employ gradient-free reinforcement learning based optimization to discover discrete prompts, especially when LLMs are accessible only via APIs (i.e., model gradients and weights are not accessible).
@@
    Transferable Prompt Tuning: Recent studies have also investigated the transferability of prompts (?), where soft prompts are first learned from one or more source tasks and then used as the prompt initialization for the target task. This is motivated by the findings (?, ?) that a good prompt initialization is crucial for prompt tuning to achieve competitive performance on the target task, compared to model tuning, especially when the model sizes of LLMs are small. Follow-up research (?) further analyzes the correlation between soft prompt transferability and the overlapping rate of activated neurons. Inspired by multi-task learning (?), MPT (?) decomposes the soft prompts for source tasks into a shared matrix and low-rank task-specific matrices, and then transfers the shared matrix to the target tasks. Additionally, studies (?, ?) have explored the transferability of soft prompts across different language models in zero/few-shot learning settings.
3.2.3 Transferability on Time Series Data
A time series is a sequence of observations collected at even intervals of time and ordered chronologically (?). Time series has been extensively applied to model non-stationary data in various high-impact domains, such as weather monitoring (?), financial forecasting (?), and healthcare (?). The key challenge in time series analysis lies in characterizing the temporal dependencies and non-stationary (i.e., rapidly changing data distribution over time) of time series data. Generally, time series transfer learning involves the following two tasks: time series forecasting (?, ?) and classification (?).
Definition 5 (Time Series Transferability for Forecasting (?))

### v2

@@ 3.2. Non-IID Transferability
The IID assumption is often violated in real scenarios, e.g., connected nodes in graphs (GCN), word occurrence in texts (LeeDS18), temporal observations in time series (VRADA), etc. To bridge this gap, non-IID transferability explores the knowledge transfer across domains, assuming that samples within each domain can be interdependent.
@@
Figure 4. Illustration of distribution shifts in characterizing graph transferability. The node distribution within the graph can be represented by $\mathrm{P}(X,G,Y)$ where $X,G,Y$ denote the input node attributes, topology structure, and output class labels, respectively. The color of nodes indicates the class labels $Y$ (blue or green).
3.2.1. Transferability on Graph Data
@@
Graph data is being generated across a variety of application domains, ranging from bioinformatics (MPNN) to e-commerce (GRADE), from protein-protein interaction prediction (GraphSAGE) to social network analysis (GIN). To capture the complex structure of graph data, graph neural networks (GNNs) (ScarselliGTHM09; DefferrardBV16) have been introduced to encode the nodes within the graphs into low-dimensional vector representations. It is shown (zhang2019graph) that there are two major learning paradigms for GNNs: spectral-based (GCN; DefferrardBV16) and spatial-based GNNs (GraphSAGE; GIN; MPNN). Recently, the transferability of spectral and spatial GNNs has been studied by exploring whether GNNs are transferable across graphs of varying sizes and topologies (WNN; GRADE). In this survey, we focus on understanding the transferability of GNNs in node-level graph learning tasks. Note that the key challenge to theoretically understand the transferability of GNNs is to measure the distribution shifts of two graphs. As illustrated in Fig. 4, the distribution shifts between source and target graphs are generally induced by the joint probabilities $\mathrm{P}_{S}(X,G,Y)$ and $\mathrm{P}_{T}(X,G,Y)$, where $X,G,Y$ denote the input node attributes, topology structure, and output class labels, respectively. Specifically, the transferability of spectral GNNs leverages the graph limits, e.g., graphon (WNN; maskey2023transferability) and graphop (LeJ23), to determine if two graphs represent the same underlying structure as the number of nodes goes to infinity. In contrast, the transferability of spatial GNNs typically relies on the empirical distribution differences of node representations in a latent embedding space learned by GNNs (GRADE).
@@
Spectral GNNs define the graph convolutions in the spectral domain using the graph Fourier transform from the perspective of graph signal processing (DefferrardBV16; GCN). Recent efforts (RuizGR21; levie2021transferability; maskey2023transferability; WNN; LeJ23) have been dedicated to understanding the transferability of spectral GNNs by answering the following question: Can spectral GNNs trained on a source graph perform well on a target graph of different sizes? This question is also known as size generalization (YehudaiFMCM21; BevilacquaZ021). The intuition behind the transferability of spectral GNNs is that if two graphs represent the same underlying phenomenon, their GNN outputs will be similar. Thus, the transferability of spectral GNNs can be derived from various aspects, including generic topological space (levie2021transferability), graphon (WNN), graphop (LeJ23), and $k$-hop ego-graph (EGI). To be more specific, (levie2019transferability; levie2021transferability) study the transferability of spectral graph filters on different discretizations of the same underlying continuous topological space. Later, graphon theory (lovasz2012large) is used to analyze the transferability of spectral GNNs. Formally, a graphon is defined by a bounded symmetric kernel and can be viewed as a graph with an uncountable number of nodes. In particular, (WNN) leverage graphon to study the asymptotic behavior of GNNs (DefferrardBV16), showing that GNNs converge to graphon neural networks (WNNs) as the number of nodes increases to infinity. This convergence implies that under mild assumptions, GNNs are transferable across graphs with performance guarantees if both graphs are drawn from the same graphon (WNN; RuizGR21; maskey2023transferability). Following this observation, recent works (CervinoRR23; KrishnagopalR23) further demonstrate the transferability of the gradients of spectral-based GNNs across graphs under similar conditions. Furthermore, using the graphop operator (backhausz2022action), (LeJ23) extend the transferability analysis of GNNs to both dense and sparse graphs. Besides, assuming that the $k$-hop ego-graphs are independent and identically drawn, (EGI) derive the transferability of a well-designed GNN based on the differences of $k$-hop ego-graph Laplacian across graphs.
@@
Spatial GNNs generally follow a recursive message-passing scheme (MPNN), where each node updates its feature vector by aggregating the message from its local neighborhood. As discussed in (GRADE; StruRW), the marginal distribution shifts $\mathrm{P}(X,G)$ between source and target graph domains can be induced by graph structure and individual node attributes (see Fig. 4(a)-(c)). Notably, three frameworks have been developed to enhance the transferability of spatial GNNs: invariant node representation (GRADE; YouCWS23; EvoluNet), structure reweighting (StruRW; Pair-Align), and graph Gaussian process (GraphGP).
@@
    Invariant Node Representation: Inspired by the domain adaptation theory (redko2019advances), it is theoretically shown (GRADE; YouCWS23) that the target error can be bounded in terms of the source error and the graph domain discrepancy. The crucial idea of invariant node representation learning is to explicitly minimize the graph domain discrepancy in a latent feature space, thereby enhancing the transferability of spatial GNNs. For example, AdaGCN (AdaGCN) and UDA-GCN (UDAGCN) leverage a domain discriminator to learn the domain-invariant node representation learned by the output layer of GNNs. Inspired by the connection between spatial GNNs and Weisfeiler-Lehman graph kernels (weisfeiler1968reduction; WLKernels), GRADE (GRADE) is proposed based on a graph subtree discrepancy measuring the subtree representation induced distribution shifts across graphs. More recently, SpecReg (YouCWS23) and A2GNN (A2GNN) further discuss the impact of spectral regularization and asymmetric model architectures on the transferability of GNNs, respectively.
@@
    Structure Reweighting: It is noticed (StruRW) that invariant node representation might lead to sub-optimal solutions under conditional structure shifts. To solve this problem, StruRW (StruRW) and Pair-Align (Pair-Align) are proposed to reweigh the edges of the source graph based on the label-oriented node connections of source and target graphs.
@@
    Graph Gaussian Process: Spatial GNNs are equivalent to graph Gaussian processes in the limit as the width of graph neural layers approaches infinity (NiuA023). Based on this observation, GraphGP (GraphGP) is derived from a graph structure-aware neural network in the limit on the layer width, in order to characterize the relationships between nodes across different graph domains. The generalization analysis of GraphGP further reveals the positive correlation between knowledge transferability and graph domain similarity.
3.2.2. Transferability on Textual Data
@@
Transfer learning has been widely studied in various natural language processing (NLP) tasks, e.g., text classification (RuderH18), question answering (WieseWN17), neural machine translation (zhao2020learning), etc. A key challenge in understanding textual transferability is the non-IID nature of words/tokens, as they might co-occur within sequences or documents. Thus, recent theoretical analyses of textual transferability often consider an alternative assumption (lotfi2024nonvacuous), i.e., sequences or documents are independently drawn from the same distribution. This assumption enables theoretically deriving the transferability and generalization of transfer learning in sequence-level and document-level NLP tasks. Take multilingual machine translation as an example, the goal is to train a single neural machine translation model to translate between multiple source and target languages (ZophK16). To achieve this, language-invariant representation learning has been introduced to align the sentence distributions of different languages within a shared latent space (arivazhagan2019missing). Nevertheless, (zhao2020learning) theoretically analyze the fundamental limits of language-invariant representation learning by deriving a lower bound (w.r.t. marginal sentence distributions from different languages) on the translation error in the many-to-many language translation setting.
@@
More recently, large language models (LLMs) have revolutionized the field of NLP (GPT4; T5; GPT3; Llama). The transferability of LLMs has been studied, as fine-tuning LLMs on downstream tasks has become the de facto learning paradigm. However, it is computationally expensive and resource-intensive to fine-tune the entire LLM model weights with billions of parameters via gradient-based optimization (BERT). To solve this problem, parameter-efficient fine-tuning (PEFT) has been investigated from the perspectives of model tuning (BitFit; LoRA; Adapters) and prompt tuning (AutoPrompt; Prefix-Tuning; PromptTuning). The goal is to adapt LLMs to various downstream tasks by adjusting as few parameters as possible.
@@
Model tuning based approaches explore model architectures or parameters of LLMs for parameter-efficient fine-tuning (AghajanyanGZ20). A variety of parameter-efficient model tuning frameworks have been proposed, including adapters (Adapters), low-rank decomposition (LoRA; Compacter), and selective masking (BitFit; DiffPruning). The key ideas behind these frameworks are illustrated in Fig. 5. In general, these approaches aim to update only a few parameters by inserting new trainable modules, adding low-rank parameter matrices, or modifying specific parameters (e.g., bias terms).
@@
    Adapters: Adapter-based approaches add new learnable modules with a small number of parameters to LLMs, e.g., maximizing the likelihood $p(y|x;\theta_{\mathrm{adapter}}\circ\theta_{\mathrm{LLM}})$ with added modules $\theta_{\mathrm{adapter}}$. Initially, inspired by visual adapter modules (RebuffiBV17), (Adapters) study the adapter-based fine-tuning mechanism in NLP tasks. This method inserts two adapters sequentially within each Transformer block (Attention): one following the self-attention layer and another after the feed-forward layer. Nevertheless, follow-up research (AdapterFusion; BapnaF19) demonstrates that inserting a single adapter after the feed-forward layer can achieve competitive performance while adding fewer parameters. Furthermore, by highlighting the connections between (model-based) adapters and (prompt-based) Prefix Tuning (Prefix-Tuning), (HeZMBN22) introduce a family of parallel adapters that directly condition the adapters at different Transformer layers on the input text. On top of adapters, AdapterDrop (AdapterDrop) and CoDA (ConditionalAdapters) improve both fine-tuning and inference efficiency by removing adapters from lower Transformer layers and querying only a small subset of input tokens against the pre-trained LLMs, respectively.
    (a) Pre-trained model with Transformer layers
    (b) Compositional parameter updates (inserted adapter modules)
    (c) Additive parameter updates (added low-rank $\Delta W$)
    (d) Selective parameter updates (selected bias term only)
    Figure 5. Illustration of parameter-efficient fine-tuning with (b) Adapters (Adapters) where new modules are inserted, (c) LoRA (LoRA) where low-rank parameter matrices are added, and (d) BitFit (BitFit) where only bias terms will be updated.
@@
    Low-rank Decomposition: Low-rank decomposition injects trainable low-rank decomposition matrices into pre-trained model parameters without changing the model architectures, e.g., maximizing the likelihood $p(y|x;\theta_{\mathrm{LLM}}+\Delta\theta_{\mathrm{LLM}})$ with low-rank $\Delta\theta_{\mathrm{LLM}}$. This line of research is motivated by the phenomenon (AghajanyanGZ20; LiFLY18) that pre-trained models tend to have a low intrinsic dimension. Here, the intrinsic dimension indicates the lowest dimensional parameter subspace in which satisfactory fine-tuned accuracy on downstream tasks can be achieved. Moreover, by assuming that the parameter change in LLMs during fine-tuning also has a low ”intrinsic rank”, LoRA (LoRA) is introduced by optimizing low-rank decomposition matrices of the parameter change. Empirically, LoRA has been further improved in various aspects, including rank selection/optimization (DyLoRA; AdaLoRA; SoRA), advanced optimizer (LoRA+; zhang2024riemannian), etc. Theoretically, the expressiveness and generalization of LoRA have also been analyzed. Notably, (zeng2024the) prove that under mild conditions regarding the rank of LoRA, it can adapt a pre-trained (or randomly initialized) model to approximate any target model of equal or smaller size. (MalladiWYCA23) find that LoRA fine-tuning is approximately equivalent to full fine-tuning in the Neural Tangent Kernel (NTK) regime (NTK), if $r\geq\Theta(\log n^{t}/\epsilon^{2})$ where $r$ is the rank of LoRA and $\epsilon$ is an approximation tolerance. (jang2024lora) show that LoRA fine-tuning has no spurious local minima in the NTK regime, if $r(r+1)>2Kn^{t}$ where $K$ is the output dimension, and $n^{t}$ is the number of training samples in the downstream target task. Furthermore, the generalization bounds of LoRA fine-tuning are theoretically derived in recent works (jang2024lora; zhu2024asymmetry).
@@
    Selective Masking: The crucial idea of selective masking is to update only a small subset of model parameters during fine-tuning, e.g., maximizing the likelihood $p(y|x;\theta_{\mathrm{LLM}}+\Delta\theta_{\mathrm{LLM}})$ with extremely sparse $\Delta\theta_{\mathrm{LLM}}$. Intuitively, it aims to find a binary mask that automatically selects a small subset of parameters for fine-tuning (ZhaoLMJS20). There are three main frameworks for learning this mask. The first one is random masking (xu2024random; Child-Tuning), where all the elements of a mask are sampled independently from a Bernoulli distribution. (xu2024random) demonstrate the effectiveness of random masking under a larger-than-expected learning rate, by theoretically building the connection between random masking and flat loss landscape. The second approach involves heuristically-motivated masks, such as bias terms (BitFit) and cross-attention layers (Gheini0M21). The third approach optimizes masks over model parameters from various perspectives, e.g., $L_{0}$-norm penalty (DiffPruning), Fisher information (FISH-Mask; Child-Tuning; FishDip), Lottery Ticket Hypothesis (LT-SFT; PlonerA24), etc.
Unlike model tuning, prompt tuning keeps the pre-trained LLMs fixed and pretends a sequence of virtual token embeddings (referred to as a trainable prompt) to the input text, e.g., maximizing the likelihood $p(y|[x,z];\theta_{\mathrm{LLM}})$ for a labeled sample $(x,y)$ where $\theta_{\mathrm{LLM}}$ denotes LLM model parameters and $z$ represents a prompt. Generally, there are three main frameworks for optimizing newly added prompts: soft (continuous) prompt tuning (Prefix-Tuning; PromptTuning), hard (discrete) prompt tuning (AutoPrompt), and transferable prompt tuning (SPoT).
@@
    Soft Prompt Tuning: The key idea of soft prompt tuning is to represent the virtual prompt as continuous-valued token embeddings. It will update only the continuous-valued embeddings of prompts $z$, either in the input embedding layer (PromptTuning; ResidualPT) or in different layers of the LLMs (Prefix-Tuning; P-Tuning; WARP; QinE21). Theoretically, (WeiXM21) studies the connection between prompt tuning and downstream tasks using an underlying latent variable generative model of text. By assuming that input texts are generated by a Hidden Markov Model (HMM), this work models the downstream task as a function of the posterior distribution of the latent variables. It is then shown that prompt tuning enhances the recovery of the ground-truth labeling function in the downstream classification task. Later, (WangC0H23) further show that a carefully constructed pre-trained Transformer can leverage prompt tuning to approximate any sequence-to-sequence function in a Lipschitz function space. They also analyze the restricted expressiveness of prompt tuning compared to model fine-tuning (e.g., LoRA (LoRA)).
@@
    Hard Prompt Tuning: Although soft prompts can be optimized via gradient-based optimization, (KhashabiLMQ0WHK22) reveal that the learned embeddings of soft prompts do not correspond to any human-readable tokens, thus lacking semantic interpretations. An alternative solution is hard prompt optimization (AutoPrompt; GrIPS), which aims to find human-readable prompts from a pre-defined vocabulary. Specifically, AutoPrompt (AutoPrompt) greedily selects the optimal token for each location in the prompt based on the gradient of the loss w.r.t. the embeddings over labeled training samples. However, the greedy search strategy can result in disfluent and unnatural prompts. To solve this problem, FluentPrompt (FluentPrompt) and PEZ (PEZ) utilize projected gradient descent optimization to update all the tokens in the prompt. The crucial idea is to project the learned continuous-valued embeddings to their nearest neighbors in a pre-defined discrete token space, and then use the mapped tokens to calculate the gradient of the loss. Besides, (choi2024hard; RLPrompt) employ gradient-free reinforcement learning based optimization to discover discrete prompts, especially when LLMs are accessible only via APIs (i.e., model gradients and weights are not accessible).
@@
    Transferable Prompt Tuning: Recent studies have also investigated the transferability of prompts (SPoT), where soft prompts are first learned from one or more source tasks and then used as the prompt initialization for the target task. This is motivated by the findings (SPoT; PPT) that a good prompt initialization is crucial for prompt tuning to achieve competitive performance on the target task, compared to model tuning, especially when the model sizes of LLMs are small. Follow-up research (SuWQCLWWLLL0SZ22) further analyzes the correlation between soft prompt transferability and the overlapping rate of activated neurons. Inspired by multi-task learning (MisraSGH16), MPT (WangPKF0K23) decomposes the soft prompts for source tasks into a shared matrix and low-rank task-specific matrices, and then transfers the shared matrix to the target tasks. Additionally, studies (SuWQCLWWLLL0SZ22; wu2024zeroshot) have explored the transferability of soft prompts across different language models in zero/few-shot learning settings.
3.2.3. Transferability on Time Series Data
A time series is a sequence of observations collected at even intervals of time and ordered chronologically (chatfield2004timeseries). Time series has been extensively applied to model non-stationary data in various high-impact domains, such as weather monitoring (Dish-TS), financial forecasting (ZhouZZLH20), and healthcare (ragab2023adatime). The key challenge in time series analysis lies in characterizing the temporal dependencies and non-stationary (i.e., rapidly changing data distribution over time) of time series data. Generally, time series transfer learning involves the following two tasks: time series forecasting (DAIN; LiuWWL22) and classification (VRADA).
Definition 3.5 (Time Series Transferability for Forecasting (DAIN)).